OpenStudy (anonymous):

given the exponential equation 2x=8, what is the logarithmic form of the equation in base 10

OpenStudy (alexwee123):

i dont see how 2x=8 is an exponential equation o.o

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm 2^x=8 }\) Use the following rule, \(\large\color{midnightblue}{ \rm Log_AB=C ~~~~~->~~~~~~~~~A^C=B }\)

OpenStudy (solomonzelman):

Or you can say \(\large\color{midnightblue}{ \rm 2^x=8 }\) \(\large\color{midnightblue}{ \rm 2^x=2^3 }\) \(\large\color{midnightblue}{ \rm x=3 }\)

geerky42 (geerky42):

You just need logarithm form of this equation, so I don't think solving for x is necessary. Just use log rule SolomonZelman showed you.

geerky42 (geerky42):

hmm, base have to be 10

OpenStudy (anonymous):

so would it be log2 8/log2 10?

OpenStudy (triciaal):

given 2x = 8 as above 8 = 2^3 log (base 2) 2x = 3 to convert to base 10 log (base 10) 2x divided by log (base 10) 2 = 3